# CLEAR: Character Unlearning in Textual and Visual Modalities

### v3

@@ 5.1 Evaluation Metrics
5.2 Implementation
For the source model, we use LLaVa model Liu et al. (2023) with ViT Dosovitskiy et al. (2021) as visual encoder and LLaMa2-7B Touvron et al. (2023) as a language model. First, we finetune it on the image captioning task using the full CLEAR, both visual and textual parts. We call this model “base”, as it contains forget and retain sets. Then, we perform the unlearning process on it. We use the same hyperparameters for each method. We evaluate the unlearned model based on our metrics from Sec. 5, using the Multi-choice VQA task for the probability score and the image captioning task for the Truth Ratio. For comparison, we also present the metrics of the “gold” model. Experiment results and metrics are shown in Tab. 1, with details provided in Appx. H.
6 Results
@@
  1. RQ1:
  2. RQ2:
  3. RQ3:
6.1 Transferability from Single Domain
To investigate RQ1, we analyse how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings. For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx. F and G. For multimodal unlearning, we use our benchmark.
6.2 Impact of Modality Selection on Unlearning
Text-only Unlearning.
Visual-only Unlearning.
Multimodal Unlearning.
@@ Appendix B Unlearning Methods
    Teacher-Student (SCRUB) Kurmanji et al. (2023) The main idea of this method is to train a student model, which is taken as a desired unlearned model from the original one, such that it will "disobey" the teacher original model on the forget set. The resulting loss of student model in this method is constructed as follows:
@@ Appendix D A sample of dataset
Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in the Table 6.
@@ Appendix F Textual-only Unlearning
For unlearning of the textual domain only, we use the TOFU benchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively. The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of 1e-5 weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter. For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout. Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures. To assess the unlearning quality, we compare the obtained unlearned model with the "gold" one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics. The full results are available in table 4.
@@ Appendix G Visual-only Unlearning
For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model. From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see section I) on which the quality of the forgetting algorithms will be tested. Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities. Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
The full results are available at table 5
@@ Appendix H Multimodal unlearning hyperparameters
Appendix I U-MIA and U-LIRA

### v4

@@ 6.1 Transferability from Single Domain
6.2 Impact of Modality Selection on Unlearning
Text-only Unlearning.
Visual-only Unlearning.
Multimodal Unlearning.
@@ 7 Conclusion
Acknowledgements
The work was supported by a Research Center for Trusted Artificial Intelligence the Ivannikov Institute for System Programming of the Russian Academy of Sciences.
@@ Appendix B Unlearning Methods
    Teacher-Student (SCRUB) Kurmanji et al. (2023) The main idea of this method is to train a student model, which is taken as a desired unlearned model from the original one, such that it will “disobey” the teacher original model on the forget set. The resulting loss of student model in this method is constructed as follows:
@@ Appendix C The process of face generation
Method Real Retain Forget Forget Quality
Gold 0.27 0.66 0.05 1.00
Original 0.27 0.65 0.40 0.97
GD 0.20 0.15 0.05 0.36
GA 0.17 0.10 0.05 0.89
IDK 0.04 0.65 0.40 0.98
KL 0.18 0.04 0.05 0.87
LLMU 0.07 0.63 0.33 0.99
NPO 0.16 0.20 0.13 0.97
Retain FT 0.25 0.66 0.32 0.98
SCRUB 0.25 0.65 0.39 0.98
Table 6: Comparison of unlearning methods on the QwenVL2 model.
@@ Appendix D A sample of dataset
Our dataset consists of 200 fictitious authors, each with 15-20 visual and 20 textual questions. We add an example of data for a single person in Tab. 7.
@@ Appendix F Textual-only Unlearning
For unlearning of the textual domain only, we use the TOFU benchmark, containing question-answer pairs of about 200 authors, 20 for each of them (4000 pairs in total), and use the splits of size 90% and 10% of the entire data for retain and forget parts respectively. The "Gold" model for the further unlearning quality evaluation is trained on the retain data only, conducting 5 epochs of training with the batch size of 4, 1 gradient accumulation step, learning rate of 1e-5 weight decay of 0.01, and also applying LoRA adapter with the rank 8, $\alpha=32$ and 0 dropout parameter. For the unlearning, we first finetune the model on the entire data split with the same hyperparameters: 5 epochs of training, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, 0 dropout coefficient. Then, unlearning methods are conducted on the forget data with the following hyperparameters: 5 epochs of unlearning, batch size of 4, 1 gradient accumulation step, learning rate of 1e-5, weight decay of 0.01, LoRA rank of 8, $\alpha=32$, zero probability dropout. Such experimental settings and hyperparameters are the same for both Llama2-7B and Mistral architectures. To assess the unlearning quality, we compare the obtained unlearned model with the “gold” one and calculate ROUGE-L on retain and forget parts, Forget Quality and Model Utility metrics. Full results are available in Tab. 4.
@@ Appendix G Visual-only Unlearning
For each of these 128 models, we run the forgetting algorithm on the forget subset of this particular model. From the resulting 128 models, we randomly select 64 target models (the remaining 64 will be used as shadow models for U-MIA and U-LIRA methods, see Appx. J) on which the quality of the forgetting algorithms will be tested. Each of the 64 target models forgets a sample $\mathcal{D}f$ of 20 personalities. Additionally, for each target model, we form a holdout set $D_{H}$ by selecting 20 personalities that were not used in the training of this model.
The full results are available in Table 5.
@@ Appendix H Multimodal unlearning hyperparameters
Appendix I Multimodal unlearning on QwenVL series
In addition to our experiments on the LLAVA series, we provide results for the QwenVL2-2B model in the table 6. We use same hyperparameters as stated in H, except that we do not use the LoRA adapters.
Appendix J U-MIA and U-LIRA